40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 80 of 329  Not logged in ELOG logo
ID Date Author Type Category Subject
  12582   Thu Oct 27 09:38:32 2016 SteveUpdatePEMmouse

We may have a mouse in the lab.  Do not leave any food scrap in trash ! Traps will be set.

 

Attachment 1: mouse.jpg
mouse.jpg
  12581   Wed Oct 26 16:06:01 2016 JohannesUpdateGeneralAutolocker maintenance

[Gautam, Johannes]

The autolocker was acting up today, Gautam traced it to EPICS channels ( namely C1:IOO-MC_LOCK_ENABLE and C1:IOO-MC_AUTOLOCK_BEAT ) served by c1iool0 not being responsive and keyed the crate. This restored it nominal operation.

  12580   Tue Oct 25 18:07:28 2016 KojiUpdateGeneralPRFPMI locked, arms loss improved

Great to hear that we have the PRG of ~16 now!

Is this 150ppm an avg loss per mirror, or per arm?

  12579   Tue Oct 25 15:56:11 2016 gautamUpdateGeneralPRFPMI locked, arms loss improved

[ericq,gautam]

Given that most of the post vent recovery tasks were done, and that the ALS noise performance looked good enough to try locking, we decided to try PRFPMI locking again last night. Here are the details:


PRM alignment, PRMI locking

  • We started by trying to find the REFL beam on the camera, the alignment biases for the 'correct' PRM alignment has changed after the vent
  • After aligning, the Oplev was way off center so that was fixed. We also had to re-center the ITMX oplev after a few failed locking attempts
  • The REFL beam was centered on all the RFPDs on the ASDC table

Post the most recent vent, where we bypass the OMC altogether, we have a lot more light now at the AS port. It has not yet been quantified how much more, but from the changes that had to be made to the loop gain for a stable loop, we estimate we have 2-3 times more power at the AS port now.


PRFPMI locking

  • We spent a while unsuccessfully trying to get the PRMI locked and reduce the carm offset on ALS control to bring the arms into the 'buzzing' state - the reason was that we forgot that it was established a couple of weeks ago that REFL165 had better MICH SNR. Once this change was made, we were readily able to reduce the carm offset to 0
  • Then we spent a few attempts trying to do blend in RF control - as mentioned in the above referenced elog, the point of failure always was trying to turn on the integrator in the CARM B path. We felt that the appearance of the CARM B IN1 signal on dataviewer was not what we are used to seeing but were unable to figure out why (as it turns out, we were locking CARM on POY11 and not REFL11 indecision, more on this later)
  • Eric found that switching the sign of the CARM B gain was the solution - we spent some time puzzling over why this should have changed, and hypothesized that perhaps we are now overcoupled, but it is more likely that this was because of the error signal mix up mentioned above...
  • We also found the DC coupling of the ITM Oplev loops to be not so reliable - perhaps this has to do with the wonky ITMY UL OSEM, more on this later. We usually turn the DC coupling on after dither aligning the arms, and in the past, it has been helpful. But we had more success last night with the DC coupling turned off rather than on.
  • Once the sign flip was figured out, we were repeatedly able to achieve locks with CARM partially on RF - we got through about 3 or 4, each was stable for just tens of seconds though. Also, we only progressed to RF on CARM on 1 attempt, the lock lasted for just a few seconds
  • Unfortunately, the mode cleaner decided to act up just about after we figured all this out, and it was pushing 4am so we decided to give up for the night.
  • The arm transmissions hit 300! We had run the transmission normalization scripts just before starting the lock so this number should be reliable (compare to ~130 in October last year). The corresponding PRG is about 16.2, which according to my Finesse models suggest we are still undercoupled, but are close to critical coupling (this needs a bit more investigation, supporting plots to follow). => Average arm loss is ~150ppm! So looks like we did some good with the vent, although of course an independent arm loss measurement has to be done...
  • Lockloss plot for one of the locks is Attachment #1

Other remarks:

  • Attachment #2 shows that the ITMY UL coil is glitchy (while the others are not). At some point last night, we turned off this sensor input to the damping servos, but for the actual locks, we turned it back on. I will do a Satellite box swap to see if this is a Sat. Box problem (which I suspect it is, the bad Sat. Boxes are piling up...)
  • Just now, eric was showing me the CM board setup in the LSC rack, because for the next lock attempts, we want to measure the CARM loop - but we found that the input to the CM board was POY and not REFL! This probably explains the sign flip mentioned above. The mix-up has been rectified
  • The MICH dither align doesn't seem to be working too well - possibly due to the fact that we have a lot more ASDC light now, this has to be investigated. But last night, we manually tweaked the BS alignment to make the dark port dark, and it seemed to work okay, although each time we aligned the PRMI on carrier, then went back to put the arms on ALS, and came back to PRMI, we would see some yaw misalignment in the AS beam...
  • I believe the SRM sat. box is still being looked at by Ben so it has not been reinstalled...
  • Eric has put together a configure script for the PRFPMI configuration which I have added to the IFO configure MEDM screen for convenience
  • For some reason, the appropriate whitening gain for POX11 and the XARM loop gain to get the XARM to lock has changed - the appropriate settings now are +30dB and 0.03 respectively. These have not been updated in some scripts, so for example, when the watch script resets the IFO configuration, it doesn't revert to these values. Just something to keep in mind for now...
Attachment 1: PRFPMIlock_25Oct2016.pdf
PRFPMIlock_25Oct2016.pdf
Attachment 2: ITMYwoes.png
ITMYwoes.png
  12578   Mon Oct 24 11:39:13 2016 gautamUpdateGeneralALS recovered

I worked on recovering ALS today. Alignments had drifted sufficiently that I had to to the alignment on the PSL table onto the green beat PDs for both arms. As things stand, both green (and IR) beats have been acquired, and the noise performance looks satisfactory (see Attachment #1), except that the X beat noise above 100Hz looks slightly high. I measured the OLTF of the X end green PDH loop (after having maximized the arm transmission, dither alignment etc, measurement done at error point with an excitation amplitude of 25mV), and adjusted the gain such that the UGF is ~10kHz (see Attachment #2).

Attachment 1: ALSOutOfLoop20161024.pdf
ALSOutOfLoop20161024.pdf
Attachment 2: XendPDHOLTF20161024.pdf
XendPDHOLTF20161024.pdf
  12577   Fri Oct 21 09:28:21 2016 SteveUpdateVACVac Normal reached

Dry pump of TP3 replaced after 9.5 months of operation.[ 45 mTorr d3 ]

The annulosses are pumped.

Valve configuration: vac normal, IFO pressure 4.5E-5 Torr [1.6E-5 Torr d3 ] on new ITcc gauge, RGA is not installed yet.

Note how fast the pressure is dropping when the vent is short.

Quote:

IFO pressure 1.7E-4 Torr on new not logged cold cathode gauge. P1 <7E-4 Torr

Valve configuration: vac.normal with anunulossess closed off.

TP3 was turned off with a failing drypump. It will be replaced tomorrow.

All time stamps are blank on the MEDM screens.

Attachment 1: VacNormal.png
VacNormal.png
  12576   Fri Oct 21 02:06:20 2016 gautamUpdateGeneralIFO recovery

The pressure on the newly installed gauge on the X arm was 6E-5 torr when I came in today evening, so I decided to start the recovery process.

  1. I first tried working at low power. I was able to lock the IMC as well as the arms. But the dither alignment didn't work so well. So I decided to go to nominal PSL power.
  2. I first changed the 2" HR mirror that is used to send all the MC REFL light to the MC REFL PD in low power operation with a 10% BS. I then roughly aligned the beam onto the PD using the tiny steering mirror. At this point, I also re-installed the ND filters on the end Transmon QPDs and also the CCD at the Y end.
  3. I then rotated the waveplate (the second one from the PSL aperture) until I maximized the power as measured just before the PSL shutter with a power meter. I then re-aligned the PMC to maximize transmission. After both these steps, we currently have 1.09W of IR light going into the IMC
  4. I then re-aligned MC REFL onto the PD (~90mW of light comes through to the PD) and maximized the DC output using an oscilloscope. I then reverted the Autolocker to the nominal version from the low power variant that has been running on megatron during the vent (although we never really used it). The autolocker worked well and I was able to lock the IMC without much trouble. I tweaked the alignment sliders for the IMC optics, but wasn't able to improve the transmission much. It is ~14600 cts right now, which is normal I think
  5. I then centered the beams onto the WFS QPDs, ran the WFSoffsets script after turning the inputs to the WFS servos off, and ran the relief script as well - I didn't try anything further with the IMC
  6. I then tried to lock the arms - I first used the green to align the test-masses. Once I was able to lock to a green 00-mode, I saw strong IR flashes and so I was able to lock the Y arm. I then ran the dither. Next, I did the same for the X arm. Even though I ran LSCoffsets before beginning work tonight, the Y arm transmission after maximization is ~5, and that for the X arm is ~2.5. I refrained from running the normalization scripts in case I am missing something here, but the mode itself is clearly visible on the cameras and is a 00-mode.
    GV edit 21Oct2016: For the Y-arm, the discrepancy was down to TRY being derived from the high gain PD as opposed to the QPD. Switching these and running the dither, TRY now maxes out at around 1.0. For TRX, the problem was that I did not install one of the ND filters - so the total ND was 1.2 rather than 1.6, which is what we were operating at and which is the ND on TRY. Both arms now have transmission ~1 after maximizing with the dither alignment...
  7. The AS spot looks nice and round on the camera, although the real check would be to do the sort of scan Yutaro and Koji did, and monitor the ASDC levels. I am leaving this task for tomorrow, along with checking the recycling cavities.
  8. Lastly, I centered the Oplevs for all the TMs

 

  12575   Thu Oct 20 15:59:52 2016 SteveUpdateVACPumpdown 80 has completed

IFO pressure 1.7E-4 Torr on new not logged cold cathode gauge. P1 <7E-4 Torr

Valve configuration: vac.normal with anunulossess closed off.

TP3 was turned off with a failing drypump. It will be replaced tomorrow.

Attachment 1: 7hPd80.png
7hPd80.png
Attachment 2: 8hPd80.png
8hPd80.png
Attachment 3: 9hPd80.png
9hPd80.png
  12574   Thu Oct 20 09:17:08 2016 SteveUpdateVACPumpdown 80 has started

Pumping again after 7 days at atmosphere.

BS, ITMY and OMC chambers were open only.

Checked: jam nuts, viewport covers and beam shutters.

Oplev servo turned off and medm screens shots taken.

New item in vacuum:  green shade 14 glass beam block at IR-input [ from the PSL ] viewport to block green reflection-scatter.

Reminder: viewport is not AR coated for green!

 

 

 

Attachment 1: readyPd80.png
readyPd80.png
Attachment 2: beforePd80.png
beforePd80.png
Attachment 3: pd80at1h.png
pd80at1h.png
Attachment 4: 7d_atm.png
7d_atm.png
  12573   Wed Oct 19 18:32:25 2016 rana, yinziUpdatePSLRefCav thermal control: heater is dead

We wanted to re-activate the Heater for the reference cavity today to use it as a testbed for PID autotuning and the new heater driver circuit that Andrew is working on for the coating thermal noise experiment.

Unfortunately, it seems that the large power supply which is used for the heater is dead.sad Or maybe I don't remember how to use it?

The AC power cord was plugged in to a power strip which seems to work for IO chassis. We also tried swapping power strip ports.

We checked the front panel fuses. The power one was 3 Ohms and the 'bias' one was 55 Ohms. We also checked that the EPICS slider did, in fact, make voltage changes at the bias control input.

Non of the front panel lights come onfrown, but I also don't remember if that is normal.

Have those lights been dead a long time? We also reconnected the heater cable at the reference cavity side.

  12572   Wed Oct 19 17:02:34 2016 ranaUpdateGeneral Viewports & coating of 2016

These old specs are not so bad. But we now want to get replacements for the TRX and TRY and PSL viewports that are R <0.1% at 532 and 1064 nm.

I don't know of any issues with keeping BK-7 as the substrate.

  12571   Wed Oct 19 16:41:55 2016 gautamUpdateGeneralHeavy doors back on

[ericq, lydia, steve, gautam]

  • We aligned the arms, and centered the in-air AS beam onto the PDs and camera
  • Misaligned the ITMs in a controlled ramp, observed ASDC level, didn't see any strange features
  • We can misalign the ITMs by +/- 100urad in yaw and not see any change in the ASDC level (i.e. no clipping). We think this is reasonable and it is unlikely that we will have to deal with such large misalignments. We also scanned a much larger range of ITM misalignments (approximately +/-1mrad), and saw no strange features in the ASDC levels as was noted in this elog - we used both the signal from the AS110 PD which had better SNR and also the AS55 PD. We take this to be a good sign, and will conduct further diagnostics once we are back at high power.
  • Opened up all light doors, checked centering on all 6 OM mirrors again, these were deemed to be satisfactory 
  • To solve the green scattering issue, we installed a 1in wide glass piece (~7inches tall) mounted on the edge of the OMC table to catch the reflection off the window (see Attachment #1) - this catches most of the ghost beams on the PSL table, there is one that remains directly above the beam which originates at the periscope in the BS/PRM chamber (see Attachment #2) but we decided to deal with this ghost on the PSL table rather than fiddle around in the vacuum and possibly make something else worse
    Link to IMG_2332.JPG
    Link to IMG_2364.JPG
  • Re-aligned arms, ran the dither, and then aligned the PRM and SRM - we saw nice round DRMI flashes on the cameras
  • Took lots of pictures in the chamber, put heavy doors back on. Test mass Oplev spots looked reasonably well centered, I re-centerd PRM and SRM spots in their aligned states, and then misaligned both
  • The window from the OMC chamber to the AS table looked clean enough to not warrant a cleaning..
  • PSL shutter is closed for now. I will check beam alignment, center Oplevs, and realign the green in the evening. Plan is to pump down first thing tomorrow morning

AS beam on OM1

Link to IMG_2337.JPG

AS beam on OM2

AS beam on OM3

AS beam on OM4

 
AS beam on OM6

I didn't manage to get a picture of the beam on OM5 because it is difficult to hold a card in front of it and simultaneously take a photo, but I did verify the centering...

It remains to update the CAD diagram to reflect the new AS beam path - there are also a number of optics/other in-vacuum pieces I noticed in the BS/PRM and OMC chambers which are not in the drawings, but I should have enough photos handy to fix this.  

Here is the link to the Picasa album with a bunch of photos from the OMC, BS/PRM and ITMY chambers prior to putting the heavy doors back on...


SRM satellite box has been removed for diagnostics by Rana. I centered the SRM Oplev prior to removing this, and I also turned off the watchdog and set the OSEM bias voltages to 0 before pulling the box out (the PIT and YAW bias values in the save files were accurate). Other Oplevs were centered after dither-aligning the arms (see Attachment #8, ignore SRM). Green was aligned to the arms in order to maximize green transmission (GTRX ~0.45, GTRY ~0.5, but transmission isn't centered on cameras).

I don't think I have missed out on any further checks, so unless anyone thinks otherwise, I think we are ready for Steve to start the pumpdown tomorrow morning.

Attachment 1: IMG_2332.JPG
IMG_2332.JPG
Attachment 2: IMG_2364.JPG
IMG_2364.JPG
Attachment 3: IMG_2337.JPG
IMG_2337.JPG
Attachment 4: IMG_2338.JPG
IMG_2338.JPG
Attachment 5: IMG_2356.JPG
IMG_2356.JPG
Attachment 6: IMG_2357.JPG
IMG_2357.JPG
Attachment 7: IMG_2335.JPG
IMG_2335.JPG
Attachment 8: Oplevs_19Oct2016.png
Oplevs_19Oct2016.png
  12570   Wed Oct 19 14:43:15 2016 SteveUpdateGeneral Viewports & coating of 2001

Tilted viewports installed in horizontal position. Atm2

Attachment 1: vacViewp2001.PDF
vacViewp2001.PDF vacViewp2001.PDF
Attachment 2: tiltedViewport.PDF
tiltedViewport.PDF
  12569   Wed Oct 19 08:28:11 2016 SteveUpdateSUSITMY_UL

Everybody is happy, except ITMY_UL or satalite box.

Gautam shows perfect form in the OMC chamber.

Attachment 1: 12hrs.png
12hrs.png
Attachment 2: vent79.jpg
vent79.jpg
  12568   Tue Oct 18 18:56:57 2016 gautamUpdateGeneralOM5 rotated to bypass OMC, green scatter is from window to PSL table

[ericq, lydia, gautam]

  • We started today by checking leveling of ITMY table, all was okay on that front after the adjustment done yesterday. Before closing up, we will have detailed pictures of the current in vacuum layout
  • We then checked centering on OMs 1 and 2 (after having dither aligned the arms), nothing had drifted significantly from yesterday and we are still well centered on both these OMs
  • We then moved to the BS/PRM chamber and checked the leveling, even though nothing was touched on this table. Like in the OMC chamber, it is difficult to check the leveling here because of layout constraints, but I verified that the table was pretty close to being level using the small (clean) spirit level in two perpendicular directions
  • Beam centering was checked on OMs 3 and 4 and verified to be okay. Clearance of beam from OM4 towards the OMC chamber was checked at two potential clipping points - near the green steering mirror and near tip-tilt 2. Clearance at both locations was deemed satisfactory so we moved onto the OMC chamber
  • We decided to go ahead and rotate OM5 to send the beam directly to OM6 and bypass the partially transmissive mirror meant to send part of the AS beam to the OMC
  • In order to accommodate the new path, I had to remove a razor beam dump on the OMC setup, and translate OM5 back a little (see Attachment #1), but we have tried to maintain ~45 degree AOI on both OMs 5 and 6
  • Beam was centered on OM6 by adjusting the position of OM5. We initially fiddled around with the pitch and yaw knobs of OM4 to try and center the beam on OM5, but it was decided that it was better just to move OM5 rather than mess around on the BS/PRM chamber and introduce potential additional scatter/clipping
  • OMC table leveling was checked and verified to not have been significantly affected by todays work
  • It was necessary to loosen the fork and rotate OM6 to extract the AS beam from the vacuum chambers onto the AP table
  • AS beam is now on the camera, and looks nice and round, no evidence of any clipping. Some centering on in air lenses and mirrors on the AP table remains to be done. We are now pretty well centered on all 6 OMs and should have more power at the AS port given that we are now getting light previously routed to the OMC out as well. A quantitative measure of how much more light we have now will have to be done after pumping down and turning the PSL power back up
  • I didn't see any evidence of back-scattered light from the window even though there were hints of this previously (sadly the same can't be said about the green). I will check once again tomorrow, but this doesn't look like a major problem at the moment

Lydia and I investigated the extra green beam situation. Here are our findings.

  1. There appears to be 3 ghost beams in addition to the main beam. These ghosts appeared when we locked the X green and Y green individually, which lead us to conclude that whatever is causing this behaviour is located downstream of the periscope on the BS/PRM chamber
    Link to greenGhosts.JPG
  2. I then went into the BS/PRM chamber and investigated the spot on the lower periscope mirror. It isn't perfectly centered, but it isn't close to clipping on any edge, and the beam leaving the upper mirror on the periscope looks clean as well (only the X-arm green was used for this, and subsequent checks). The periscope mirror looks a bit dusty and scatters rather a lot which isn't ideal...
    Link to IMG_3322.JPG
  3. There are two steering mirrors on the IMC table which we do not have access to this vent. But I looked at the beam coming into the OMC chamber and it looks fine, no ghosts are visible when letting the main beam pass through a hole in one of our large clean IR viewing cards - and the angular separation of these ghosts seen on the PSL table suggests that we would see these ghosts if they exist prior to the OMC chamber on the card...
  4. The beam hits the final steering mirror which sends it out onto the PSL table on the OMC chamber cleanly - the spot leaving the mirror looks clean. However, there are two reflections from the two surfaces of the window that come back into the OMC chamber. Space constraints did not permit me to check what surfaces these scatter off and make it back out to the PSL table as ghosts, but this can be checked again tomorrow.
    Link to IMG_3326.JPG

I can't think of an easy fix for this - the layout on the OMC chamber is pretty crowded, and potential places to install a beam dump are close to the AS and IMC REFL beam paths (see Attachment #1). Perhaps Steve can suggest the best, least invasive way to do this. I will also try and nail down more accurately the origin of these spots tomorrow.


Light doors are back on for the night. I re-ran the dithers, and centered the oplevs for all the test-masses + BS. I am leaving the PSL shutter closed for the night

 

Attachment 1: OMCchamber.pdf
OMCchamber.pdf
Attachment 2: greenGhosts.JPG
greenGhosts.JPG
Attachment 3: IMG_3322.JPG
IMG_3322.JPG
Attachment 4: IMG_3326.JPG
IMG_3326.JPG
  12567   Tue Oct 18 17:11:42 2016 YinziUpdateGreen LockingMore serial port troubleshooting

I connected to the serial port using screen (through Terminal) and using Arduino's serial monitor and basically received the same strings that were received through python, so it's not a python issue. Checked the other TC 200 module and was also receiving nonsense, but it was all question marks instead of mostly K's and ['s.

This rules out a few possible reasons for the weird data. Next steps are to set up and configure the Raspberry Pi (which has been interfaced before) and see if the problem continues.

  12566   Mon Oct 17 22:45:16 2016 gautamUpdateGeneralAS beam centered on all OMs

[ericq, lydia, gautam]

IMC realignment, Arm dither alignment

  • We started today by re-locking the PMC (required a c1psl restart), re-locking the IMC and then locking the arms
  • While trying to dither align the arms, I could only get the Y arm transmission to a maximum of ~0.09, while we are more used to something like 0.3 when the arm is well aligned this vent
  • As it turns out, Y arm was probably locked to an HOM, as a result of some minor drift in the ITMY optical table leveling due to the SOS tower aperture being left in over the weekend

ITMY chamber

  • We then resolved to start at the ITMY chamber, and re-confirm that the beam is indeed centered on the SRM by means of the above-mentioned aperture
  • Initially, there was considerable yaw misalignment on the aperture, probably due to the table level drifting because of the additional weight of the aperture
  • As soon as I removed the aperture, eric was able to re-dither-align the arms and their transmission went back up to the usual level of ~0.3 we are used to this vent
  • We quickly re-inserted the aperture and confirmed that the beam was indeed centered on the SRM
  • Then we removed the aperture from the chamber and set about inspecting the beam position on OM1
  • While the beam position wasn't terribly bad, we reasoned that we may as well do as good a job as we can now - so OM1 was moved ~0.5 in such that the beam through the SRM is now well centered on OM1 (see Attachment #1 for a CAD drawing of the ITMY table layout and the direction in which OM1 was moved)
  • Naturally this affected the beam position on OM2 - I re-centered the beam on OM2 by first coarsely rotating OM1 about the post it is mounted on, and then with the knobs on the mount. The beam is now well centered on OM2
  • We then went about checking the table leveling and found that the leveling had drifted substantially - I re-levelled the table by moving some of the weights around, but this has to be re-checked before closing up... 

BS/PRM chamber

  • The beam from OM2 was easily located in the BS/PRM chamber - it required minor yaw adjustment on OM2 to center the beam on OM3
  • Once the beam was centered on OM3, minor pitch and yaw adjustments on the OM3 mount were required to center the beam on OM4
  • The beam path from OM3 to OM4, and OM4 to the edge of the BS/PRM chamber towards the OMC chamber was checked. There is now good clearance (>2 beam diameters) between the beam from OM4 to the OMC chamber, and the green steering mirror in the path, which was one of the prime clipping candidates identified on Friday

OMC chamber

  • First, the beam was centered on OM5 by minor tweaking of the pitch and yaw knobs on OM4 (see Attachment #2)
  • Next, we set about removing the unused mirror just prior to the window on the AP table (see Attachment #3). PSL shutter was closed for this stage of work, in order to minimize the chance of staring directly into the input beam!
  • Unfortunately, we neglected checking the table leveling prior to removing the optic. A check after removing the optic suggested that the table wasn't level - this isn't so easy to check as the table is really crowded, and we can only really check near the edges of the table (see Attachment #3). But placing the level near the edge introduces an unknown amount of additional tilt due to its weight. We tried to minimize these effects by using the small spirit level, which confirmed that the table was indeed misaligned
  • To mitigate this, we placed a rectangular weight (clean) around the region where the removed mirror used to sit (see Attachment #3)Approximately half the block extends over the edge of the table, but it is bolted down. The leveling still isn't perfect - but we don't want to be too invasive on this table (see next bullet point). Since there are no suspended optics on this table, I think the leveling isn't as critical as on the other tables. We will take another pass at this tomorrow but I think we are in a good enough state right now. 
  • All this must have bumped the table quite a bit, because when we attempted re-locking the IMC, we noticed substantial misalignment. We should of course have anticipated this because the mirror launching the input beam into the IMC, and also MMT2 launching the beam into the arms, sits on this table! After exploring the alignment space of the IMC for a while, eric was able to re-lock the IMC and recover nominal transmission levels of ~1200 counts. 
  • We then re-locked the arms (needed some tip-tilt tweaking) and ran the dither again, setting us up for the final alignment onto OM6
  • OM5 pitch and yaw knobs were used to center the beam on OM6 - the resulting beam spot on OMPO-OMMTSM and OM6 are shown in Attachment #4 and Attachment #5 respectively. The centering on OMPO-OMMTSM isn't spectacular, but I wanted to avoid moving this optic if possible. Moreover, we don't really need the beam to follow this path (see last bullet in this section)
  • Beam path in the OMC chamber (OM5 --> OMPO-OMMTSM --> OM6 --> window was checked and no significant danger of clipping was found
  • Beam makes it cleanly through the window onto the AP table. We tweaked the pitch and yaw knobs on OM6 to center the beam on the first in-air pick off mirror steering the AS beam on the AP table. The beam is now visible on the camera, and looks clean, no hint of clipping
  • As a check, I wondered where the beam into the OMC is actually going. Turns out that as things stand, it is hitting the copper housing (see Attachment #6, it's had to get a good shot because of the crowded table...). While this isn't critical, perhaps we can avoid this extra scatter by dumping this beam?
  • Alternatively, we could just bypass OMPO-OMMTSM altogether - so rotate OM5 in-situ such that we steer the beam directly onto OM6. This way, we avoid throwing away half (?) the light in the AS beam. If this is the direction we want to take, it should be easy enough to make the change tomorrow

In summary...

  • AS beam has been centered on all steering optics (OM1 through OM6)
  • Table leveling has been checked on ITMY and OMC chambers - this will be re-checked prior to closing up
  • Green-scatter issue has to be investigated, should be fairly quick..
  • In the interest of neatness, we may want to install a couple of beam dumps - one to catch the back-reflection off the window in the OMC chamber, and the other for the beam going to the OMC (unless we decide to swivel OM5 and bypass the OMC section altogether, in which case the latter is superfluous)

C1SUSAUX re-booting

  • Not really related to this work, but we couldn't run the MC relief script due to c1susaux being unresponsive
  • I re-started c1susaux (taking care to follow the instructions in this elog to avoid getting ITMX stuck)
  • Afterwards, I was able to re-lock the IMC, recover nominal transmission of ~1200 counts. I then ran the MC relief servo
  • All shutters have been closed for the night
Attachment 1: OM1Moved.pdf
OM1Moved.pdf
Attachment 2: IMG_3304.JPG
IMG_3304.JPG
Attachment 3: OMCchamber.pdf
OMCchamber.pdf
Attachment 4: IMG_3292.JPG
IMG_3292.JPG
Attachment 5: IMG_3307.JPG
IMG_3307.JPG
Attachment 6: IMG_3297.JPG
IMG_3297.JPG
  12565   Mon Oct 17 14:48:12 2016 SteveUpdatesafetysafety training

Ashley Fowler  "high shool" student received basic 40m safety training and Lydia is her guarding angle.

 

  12564   Fri Oct 14 19:59:09 2016 YinziUpdateGreen LockingContinuing work with the TC 200

Oct. 15, 2016

Another attempt (following elog 8755) to extract the oven transfer function from time series data using Matlab’s system identification functionalities.

The same time series data from elog 8755 was used in Matlab’s system identification toolbox to try to find a transfer function model of the system.

From elog 8755: H(s) is known from current PID gains: H(s) = 250 + 60/s +25s, and from the approximation G(s)=K/(1+Ts), we can expect the transfer function of the system to have 3 poles and 2 zeros.

I tried fitting a continuous-time and a discrete time transfer function with 3 poles and 2 zeros, as well as using the "quick start" option. Trying to fit a discrete time transfer function model with 3 poles and 2 zeros gave the least inaccurate results, but it’s still really far off (13.4% fit to the data).

Ideas:

1. Obtain more time domain data with some modulation of the input signal (also gives a way to characterize nonlinearities like passive cooling). This can be done with some minor modifications to the existing code on the raspberry pi. This should hopefully lead to a better system ID.

2. Try iterative tuning approach (sample gains above and below current gains?) so that a tune can be obtained without having to characterize the exact behavior of the heater.

Oct. 16, 2016

-Found the raspberry pi but it didn’t have an SD card

-Modified code to run directly on a computer connected to the TC 200. Communication seems to be happening, but a UnicodeDecodeError is thrown saying that the received data can’t be decoded.

-Some troubleshooting: tried utf-8 and utf-16 but neither worked. The raw data coming in is just strings of K’s, [‘s, and ?’s

-Will investigate possible reasons (update to Mac OS or a difference in Python version?), but it might be easier to just find an SD card for the raspberry pi which is known to work. In the meantime, modify code to obtain more time series data with variable input signals.

  12563   Fri Oct 14 18:33:55 2016 gautamUpdateGeneralAS clipping investigations

[steve,ericq,gautam]

In the afternoon, we took the heavy door off the OMC chamber as well, such that we could trace the AS beam all the way out to the AP table. 

In summary, we determined the following today:

  1. Beam is centered on SRM, as judged by placing the SOS iris on the tower
     
  2. Beam is a little off on OM1 in yaw, but still >2 beam diameters away from the edge of the steering optic, pitch is pretty good
  3. Beam is okay on OM2 
  4. Beam is okay on OM3 - but beam from OM3 to OM4 is perilously close to clipping on the green steering mirror between these two steering optics (see CAD drawing). We think this is where whatever effect of the SR2 hysteresis shows up first.
  5. Beam is a little low and a little to the left on OM4 (the first PZTJena mirror)
  6. Beam is well clear of other optics in the BS PRM chamber on the way from OM4 to OM5 in the OMC chamber
  7. Beam is a little low and a little to the left of OM5 in the OMC chamber. This is the second PZTJena mirror. We are approximately 1 beam diameter away from clipping on this 1" optic
    Link to IMG_2289.JPG
  8. Beam is off center on OMPO-OMMTSM partially transmissive optic, but because this is a 2" optic, the room for error is much more
    Link to IMG_2294.JPG
  9. Beam is well clear of optics on OMC table on the way from OMPO-OMMTSM to OM6, the final steering mirror bringing the AS beam out onto the table
  10. Beam is low and to the left on OM6. It is pretty bad here, we are < 1 beam diameter away from clipping on this optic, this along with the near miss on the BS/PRM chamber are the two most precarious positions we noticed today, consistent with the hypothesis in this elog that there could be multiple in vacuum clipping points
    Link to IMG_2306.JPG
  11. Beam clears the mirror just before the window pretty confortably (see photo, CAD drawing). But this mirror is not being used for anything useful at the moment. More importantly, there is some reflection off the window back onto this mirror frame which is then scattering and creating some ghost beams, so this could explain the anomalous ASDC behaviour Koji and Yutaro saw. In any case, I would favour removing this mirror since it is serving no purpose at the moment.
    Link to IMG_2310.JPG

Attachment #5 is extracted from the 40m CAD drawing which was last updated in 2012. It shows the beam path for the output beam from the BS all the way to the table (you may need to zoom in to see some labels. The drawing may not be accurate for the OMC chamber but it does show all the relevant optics approximately in their current positions.

EQ will put up photos from the ITMY and BS/PRM chambers.

Plan for Monday: Reconfirm all the findings from today immediately after running the dither alignment so that we can be sure that the ITMs are well-aligned. Then start at OM1 and steer the beam out of the chambers, centering the beam as best as possible given other constraints on all the optics sequentially. All shutters are closed for the weekend, though I left the SOS iris in the chamber...

Here is the link to the Picasa album with a bunch of photos from the OMC chamber prior to us making any changes inside it - there are also some photos in there of the AS beam path inside the OMC chamber...

Attachment 1: IMG_2289.JPG
IMG_2289.JPG
Attachment 2: IMG_2294.JPG
IMG_2294.JPG
Attachment 3: IMG_2306.JPG
IMG_2306.JPG
Attachment 4: IMG_2310.JPG
IMG_2310.JPG
Attachment 5: ASBeamClipping.pdf
ASBeamClipping.pdf
  12562   Fri Oct 14 15:47:00 2016 ranaUpdateTreasurefilters + clip

I say just fix the clipping. Don't worry about the PRM OSEM filters. We can do that next time when we put in the ITM baffles. No need for them on this round.

  12561   Fri Oct 14 10:31:53 2016 gautamUpdateGeneraldoors are off ITMY and BS/PRM chambers

[steve,ericq,gautam]

We re-checked IMC locking, arm alignments (we were able to lock and dither align both arms today, and also made the michelson spot look reasonable on the camera) and made sure that the AS and REFL spots were in the camera ballpark. We then proceeded to remove the heavy doors off ITMY and BS/PRM chambers. We also quickly made sure that it is possible to remove the side door of the OMC chamber with the current crane configuration, but have left it on for now.

The hunt for clipping now begins.

  12560   Thu Oct 13 19:28:14 2016 gautamUpdateGeneralin-air alignment

I did the following today to prepare for taking the doors off tomorrow.

  • Locked MC at low power
    • low power autolocker used during the last vent isn't working so well now
    • so I manually locked the IMC - locks are holding for ~30 mins and MC transmission was maximized by tweaking MC1 and MC2 alignments. The transmission is now ~1150 which is what I remember it being from the last vent
    • I had to restart c1aux to run LSCoffsets
  • Aligned arms to green using bias sliders on IFO align
    • X green transmission is ~0.4 and Y green transmission is ~0.5 which is what I remember it being before this vent
  • Removed ND filters from end Transmon QPDs since there is so little light now
  • Locked Y arm, ran the dither
    • Some tip tilt beam walking was required before any flashes were seen
    • I had to tweak the LSC gain for this to work
    • TRY is ~0.3 - in the previous vent, in air low power locking yielded TRY of ~0.6 but the 50-50 BS that splits the light between the high gain PD and the QPD was removed back then so these numbers are consistent
  • Tried locking X arm
    • For some reason, I can't get the triggering to work well - the trigger in monitor channel (LSC-XARM_TRIG_IN) and LSC-TRX_OUT_DQ are not the same, should they not be?
    • Tried using both QPD and high gain PD to lock, no luck. I also checked the error signal for DC offsets and that the demod phase was okay
    • In any case, there are TRX flashes of ~0.3 as well, this plus the reasonable green transmission makes me think the X arm alignment is alright
  • All the oplev spots are on their QPDs in the +/- 100 range. I didn't bother centering them for now

I am leaving all shutters closed overnight.

So I think we are ready to take the doors off at 8am tomorrow morning, unless anyone thinks there are any further checks to be done first.


Vent objectives:

  1. Fix AS beam clipping issues (elog1, elog2)
  2. Look into the green scatter situation (elog)

Should we look to do anything else now? One thing that comes to mind is should we install ITM baffles? Or would this be more invasive than necessary for this vent?


Steve reported to me that he was unable to ssh into the control room machines from the laptops at the Xend and near the vacuum rack. The problem was with pianosa being frozen up. I did a manual reboot of pianosa and was able to ssh into it from both laptops just now.

  12559   Thu Oct 13 16:48:34 2016 SteveUpdateVACvent 79 is completed

IFO is at atmosphere. The MC can be locked in air now.

The doors will be coming off tomorrow 8am sharp.

Do we want to install the ITM baffles?

What about the found OSEM filters?

 

Attachment 1: vent_79.png
vent_79.png
  12558   Thu Oct 13 14:49:57 2016 KojiConfigurationPEMXLR(F)-XLR(M) cable took from the fibox to the Blue microphone

[Gautam Koji]

XLR(F)-XLR(M) cable for the blue microphone is missing. Steve ordered one.

We found one in the fibox setup. As we don't use it during the vent, we use this cable for the microphone.
Once we get the new one, it will go to the fibox setup.

 

  12557   Thu Oct 13 09:30:38 2016 SteveUpdateGeneralvent 79 is happening
Quote:

I have completed the following non-Steve portions of the pre-vent checklist [wiki-40m.ligo.caltech.edu]

  • Center all oplevs, transom QPDs (and IPPOS + IPANG if they are set up)
  • Align the arm cavities for IR and align the green lasers to the arms.
  • Update the SUS Driftmon values
  • Reconcile all SDF differences
  • Reduce input power to no more than 100mW (measured at the PSL shutter) by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  • Close shutter of PSL-IR and green shutters at the ends

All shutters are closed. Ready for Steve to check nuts and begin venting!

- checked all jam nuts

- checked all viewports are covered

- turned oplev servos off

- took pictures of medm screens: sus summs, aligned oplev centering, IFO& MC alignment biases and vac configuration

- checked particle counts

- checked crane operational safety

- closed V1, VM1, annuloses

- opened VV1  and vented with Airgas brand, Industrial Grade Nitrogen [ 99.99% ] to 25 Torr

- switched over to Airgas brand compressed air, Alphagas " AI UZ300 "  with Total Hydro Carbon 0.1 PPM

Attachment 1: vent79precond.png
vent79precond.png
Attachment 2: venting.png
venting.png
  12556   Thu Oct 13 00:23:54 2016 ericqUpdateVACPre-vent prep

I have completed the following non-Steve portions of the pre-vent checklist [wiki-40m.ligo.caltech.edu]

  • Center all oplevs, transom QPDs (and IPPOS + IPANG if they are set up)
  • Align the arm cavities for IR and align the green lasers to the arms.
  • Update the SUS Driftmon values
  • Reconcile all SDF differences
  • Reduce input power to no more than 100mW (measured at the PSL shutter) by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  • Close shutter of PSL-IR and green shutters at the ends

All shutters are closed. Ready for Steve to check nuts and begin venting!

  12555   Wed Oct 12 22:51:26 2016 KojiUpdateGeneralClipping could not be improved by the PZTs - NEED ANOTHER VENT

[Gautam Koji]

We engaged the HV driver to the output port PZTs, hoping to mitigate the AS port clipping. Basically, the range of the PZT is not enough to make the beam look clean. Also, our observation suggested there are possible multiple clipping in the chamber. We need another vent to make the things clearly right. Eric came in the lab and preparing the IFO for it.


1. Before the test, the test masses have been aligned with the dither servo.

2. We looked at the beam shape on the AS camera with a single bounce beam. We confirmed that the beam is hard-clipped at the upper and left sides of the beam on the video display. This clipping is not happening outside of the chamber.

3. We brought an HV power supply to the short OMC rack. There is a power supply cable with two spades. The red and black wires are +150V and GND respectively.

4. The voltage of +/-10V was applied on each of the four PZT drive inputs. We found that the motion of the beam on the camera is tiny and in any case, we could not improve the beam shape.

5. We wondered that if we are observing ANY improvement of the clipping. For this purpose, we aligned AS110 sensor every time we gave the misalignment with the PZTs. Basically, we are at the alignment to have the best power we can get. We thought this was weird.

6. Then we moved the AS port spot with the ITMX. We could clearly make the spot more round. However, this reduced the power at the AS port reduced by ~15%. When the beam was further clipped, the power went down again. Basically, the initial alignment gave us the max power we could get. As the max power was given with the clipped beam, we get confused and feel safer to check the situation with the chambers open.

During this investigation, we moved the AS port opitcs and the AS camera. So they are not too precise reference of the alignment. The PZT HV setup has been removed.

  12554   Wed Oct 12 18:09:25 2016 LydiaUpdateGeneralAcromag Progress

[Lydia, Johannes]

Johannes acquired a crate to contain the Acromag setup and wiring, and installed a rail along the bottom panel so that the ADC units will be oriented vertically with the ehternet ports facing up. We briefly talkes about what the layout should be, and are thinking of using 2 rails, one for ADCs and one for DACs. We want to design a generic front panel to accept 25 pin D-Sub inputs and maybe also BNCs, which we can use for all the Acromag crates. 

I got the epics session for the acromag to run on c1iscex and was able to access the channel values using caget on donatella. However, I get the following warning: 

cas warning: Using dynamically assigned TCP port 48154,
cas warning: but now two or more servers share the same UDP port.
cas warning: Depending on your IP kernel this server may not be
cas warning: reachable with UDP unicast (a host's IP in EPICS_CA_ADDR_LIST)

 

It seems like there might be a way to assign a port for each unit, if this is a problem. 

Also, c1iscex doens't have tmux; what's the best way to run the modbusApp and then detach? Right now I just left an epics session running in an open terminal. 

Plans:

  • Deisgn crate connections and interior layout. Set up front panel to accept desired connections. 
  • Set up the crate with the Acromag XT1221 reading the diagnostic info from the X end NPRO in the X end rack.
  • Figure out how many of each type we need to replace c1auxex functionality, and order them.
  • Generate appropriate EPICS db files for acromag based on slow machine channels.
  • Add necessary units to X end Acromag crate and read in the same inputs as c1auxex. 
  • Set up everything else to look for c1auxex channels from Acromag instead. (Not sure about nuances of this step: should we name the channels something different at first? How to find everything that relies on c1auxex? Must be careful with SUS channel connections.)
  •  Determine number of units needed to replace all slow machines, and order thm. Likewise assemble as many crates as necessary with the right connections. 
  • Once we are confident that the replacement is complete and fully functional, disconnect c1auxex and repeat process for other slow machines. 
  12553   Wed Oct 12 15:01:22 2016 steveUpdateSUSSOS ITM baffles plates are ready

The two 40 mm apeture baffles at the ends were replaced by 50 mm one. ITM baffles with 50 mm apeture are baked ready for installation.

Quote:

 Green welding glass 7" x 9"   shade #14 with 40 mm hole and mounting fixtures are ready to reduce scatter light on SOS

PEEK 450CA shims and U-shaped clips  will keep these plates damped.

 

 

Attachment 1: baffle7x9_1.5.jpg
baffle7x9_1.5.jpg
Attachment 2: baffle_holder.jpg
baffle_holder.jpg
Attachment 3: baffle_top_view.jpg
baffle_top_view.jpg
  12552   Wed Oct 12 13:34:28 2016 gautamUpdateSUSPRM LR problematic again

Looks like what were PRM problems are now seen in the SRM channels, while PRM itself seems well behaved. This supports the hypothesis that the satellite box is problematic, rather than any in-vacuum shenanigans.

Eric noted in this elog that when this problem was first noticed, switching Satellite boxes didn't seem to fix the problem. I think that the original problem was that the Al foil shorted the contacts on the back of the OSEM. Presumably, running the current driver with (close to) 0 load over 2 months damaged that part of the Satellite box circuitry, which lead to the subsequent observations of glitchy behaviour after the pumpdown. Which begs the question - what is the quick fix? Do we try swapping out the LM6321 in the LR LED current driver stage? 

GV Edit Nov 2 2016: According to Rana, the load of the high speed current buffer LM6321 is 20 ohms (13 from the coil, and 7 from the wires between the Sat. Box and the coil). So, while the Al foil was shorting the coil, the buffer would still have seen at least 7 ohms of load resistance, not quite a short circuit. Moreover, the schematic suggests that that the kind of overvoltage protection scheme suggested in page 6 on the LM6321 datasheet has been employed. So it is becoming harder to believe that the problem lies with the output buffer. In any case, we have procured 20 of these discontinued ICs for debugging should we need them, and Steve is looking to buy some more. Ben Abbot will come by later in the afternoon to try and help us debug.

  12551   Tue Oct 11 13:30:49 2016 gautamUpdateSUSPRM LR problematic again

Perhaps the problem is electrical? The attached plot shows a downward trend for the LR sensor output over the past 20 days that is not visible in any of the other 4 sensor signals. The Al foil was shorting the electrical contacts for nearly 2 months, so perhaps some part of the driver circuit needs to be replaced? If so a Satellite Box swap should tell us more, I will switch the PRM and SRM satellite boxes. It could also be a dying LED on the OSEM itself I suppose. If we are accessing the chamber, we should come up with a more robust insulating cap solution for the OSEMs rather than this hacky Al foil + kapton arrangement.

The PRM and SRM Satellite boxes have been switched for the time being. I had to adjust some of the damping loop gains for both PRM and SRM and also the PRM input matrix to achieve stable damping as the PRM Satellite box has a Side sensor which reads out 0-10V as opposed to the 0-2V that is usually the case. Furthermore, the output of the LR sensor going into the input matrix has been turned off.

 

  12550   Tue Oct 11 10:38:51 2016 SteveUpdateSUS wire standoffs update

100 Sapphire prizms arrived.

 

Attachment 1: A028_-_20161010_170319.jpg
A028_-_20161010_170319.jpg
Attachment 2: A027_-_20161010_170202.jpg
A027_-_20161010_170202.jpg
  12549   Tue Oct 11 10:15:04 2016 SteveUpdateSUSPRM LR problematic again

It started here

 

Attachment 1: PRMalfoiled.png
PRMalfoiled.png
  12548   Tue Oct 11 08:09:46 2016 Max IsiUpdateDMFsummar pages dead again

Summary pages will be unavailable today due to LDAS server maintenance. This is unrelated to the issue that Rana reported.

Quote:

I've re-submitted the Condor job; pages should be back within the hour.

Quote:

Been non-functional for 3 weeks. Anyone else notice this? images missing since ~Sep 21.

 

 

  12547   Tue Oct 11 02:48:43 2016 ericqUpdateLSCRevival Attempts

Still no luck relocking, but got a little further. I disabled the output of the problematic PRM OSEM, it seems to work ok. Looking at the sensing of the PRMI with the arms held off, REFL165 has better MICH SNR due to its larger seperation in demod angle. So, I tried the slightly odd arrangement of 33I for PRCL and 165Q for MICH. This can indefinitely hold through the buzzing resonance. However, I haven't been able to find the sweet spot for turning on the CARM_B (CM_SLOW) integrator, which is neccesary for turning up the AO and overall CARM gain. This is a familiar problem, usually solved by looking at the value far from resonance on either side, and taking the midpoint as the filter module offset, but this didn't work tonight. Tried different gains and signs to no avail.

  12546   Tue Oct 11 00:43:50 2016 ericqUpdateSUSPRM LR problematic again

Tonight, and during last week's locking, we noticed something intermittently kicking the PRM. I've determined that PRM's LR OSEM is problematic again. The signal is coming in and out, which kicks the OSEM damping loops. I've had the watchdog tripped for a little bit, and here's the last ten minutes of the free swinging OSEM signal:

Here's the hour trend of the PRM OSEMS over the last 7 days a plot of just LR since the fix on the 9th of September.

It looks like it started misbehaving again on the evening of the 5th, which was right when we were trying to lock... Did we somehow jostle the suspension hard enough to knock the foil cap back into a bad spot?

  12545   Mon Oct 10 18:34:52 2016 gautamUpdateGeneralPZT OM Mirrors

I did a quick survey of the drive electronics for the PZT OM mirrors today. The hope is that we can correct for the clipping observed in the AS beam by using OM4 (in the BS/PRM chamber) and OM5 (in the OMC chamber).

Here is a summary of my findings.

  • Schematic for (what I assume is) the driver unit (located in the short electronics rack by the OMC chamber/AS table) can be found here
  • This is not hooked up to any HV power supply. There is a (short) cable on the back that is labelled '150V' but it isn't connected to anything. There are a bunch of 150V KEPCO power supplies in 1X1, looks like we will have to lay out some cable to power the unit
  • The driver is also not connected to any fast front end machine or slow machine - according to the schematic, we can use J4, which is a Dsub 9 connector on the front panel, to supply drive signals to the two PZTs X and Y axes. Presumably, we can use this + some function generator/DC power supply to drive the PZTs. I have fashioned a cable using a Dsub9 connector and some BNC connectors for this purpose.

I hope these have the correct in-vacuum connections. We also have to hope that the clipping is downstream of OM4 for us to be able to do anything about it using the PZT mirrors. 

  12544   Mon Oct 10 17:42:47 2016 Max IsiUpdateDMFsummar pages dead again

I've re-submitted the Condor job; pages should be back within the hour.

Quote:

Been non-functional for 3 weeks. Anyone else notice this? images missing since ~Sep 21.

 

  12543   Mon Oct 10 17:27:29 2016 ranaUpdateDMFsummar pages dead again

Been non-functional for 3 weeks. Anyone else notice this? images missing since ~Sep 21.

  12542   Mon Oct 10 11:48:05 2016 gautamUpdateCDSPower-cycled c1susaux, realigned PMC, spots centered on WFS1 and WFS2

[Koji, Gautam]

We did the following today morning:

  1. I re-aligned the PMC - transmission level on the scope on the PSL table is now ~0.72V which is around what I remember it being
  2. The spot had fallen off WFS 2 - so we froze the output of the MC WFS servo, and turned the servo off. Then we went to the table to re-center the spot on the WFS. The alignment had drifted quite a bit on WFS2, and so we had to change the scale on the grid on the MEDM screen to +/-10 (from +/- 1) to find the spot and re-center it using the steering mirror immediately before the WFS. It would appear that the dark offsets are different on WFS1 and WFS2, so the "SUM" reads ~2.5 on WFS1 and ~0.3 on WFS2 when the spots are well centered
  3. Coming back to the control room, we ran the WFSoffsets script and turned on the WFS servo again. Trying to run the relief servo, we were confronted by an error message that c1susaux needed to be power cycled (again). This is of course the slow machine that the ITMX suspension is controlled by, and in the past, power cycling c1susaux has resulted in the optic getting stuck. An approach that seems to work (without getting ITMX stuck)  is to do the following:
    • Save the alignment of the optic, turn off Oplev servo
    • Move the bias sliders on IFO align to (0,0) slowly
    • Turn the watchdog for ITMX off
    • Unplug the cables running from the satellite box to the vacuum feedthrough
    • Power cycle the slow machine. Be aware that when the machine comes back on, the offset sliders are reset to the value in the saved file! So before plugging the cables back in, it would be advisable to set these to (0,0) again, to avoid kicking the optic while plugging the cables back in
    • Plug in the cables, restore alignment and Oplev servos, check that the optic isn't stuck
  4. Y green beat touch up - I tweaked the alignment of the first mirror steering the PSL green (after the beam splitter to divide PSL green for X and Y beats) to maximize the beat amplitude on a fast scope. Doing so increased the beat amplitude on the scope from about 20mVpp to ~35mVpp. A detailed power budget for the green beats is yet to be done

It is unfortunate we have to do this dance each time c1susaux has to be restarted, but I guess it is preferable to repeated unsticking of the optic, which presumably applies considerable shear force on the magnets...


After Wednesday's locking effort, Eric had set the IFO to the PRMI configuration, so that we could collect some training data for the PRC angular feedforward filters and see if the filter has changed since it was last updated. We should have plenty of usable data, so I have restored the arms now.

  12541   Mon Oct 10 09:31:25 2016 SteveUpdateSUSPRM damping restored

Local earth quake 3.7 mag  trips PRM

ETMY_UL glitch

What about the MC?

 

Attachment 1: 3.7mLomaLinda.png
3.7mLomaLinda.png
Attachment 2: ETMY_UL_glitch.png
ETMY_UL_glitch.png
Attachment 3: MC_glitcing?.png
MC_glitcing?.png
  12540   Fri Oct 7 20:56:15 2016 KojiUpdateSUSOutput matrix diagonalization

I wanted to see what is the reason to have such large coupling between pitch and yaw motions.

The first test was to check orthogonality of the bias sliders. It was done by monitoring the suspension motion using the green beam.
The Y arm cavity was aligned to the green. The damping of ITMY was all turned off except for SD.
Then ITMY was misaligned by the bias sliders. The ITMY face CCD view shows that the beam is reasonably orthogonally responding to the pitch and yaw sliders.
I also confirmed that the OPLEV signals also showed reasonablly orthogonal responce to the pitch and yaw misalignment.

=> My intuition was that the coils (including the gain balance) are OK for a first approximation.

Then, I started to excite the resonant modes. I agree that it is difficult to excite a pure picth motion with the resonance.


So I wanted to see how the mixing is frequency dependent.

The transfer functions between ITMY_ASCPIT/YAW_EXC to ITMY_OPLEV_PERROR/YERROR were measured.

The attached PDFs basically shows that the transfer functions are basically orthogonal (i.e. pitch exc goes to pitch, yaw exc goes to yaw) except at the resonant frequency.

I think the problem is that the two modes are almost degenerate but not completely. This elog shows that the resonant freq of the ITMY modes are particularly close compared to the other suspensions.
If they are completely degenerate, the motion just obeys our excitation. However, they are slightly split. Therefore, we suffer from the coupled modes of P and Y at the resonant freq.
However, the mirror motion obeys the exitation at the off resonance as these two modes are similar enough.

This means that the problem exists only at the resonant frequencies. If the damping servos have 1/f slope around the resonant freqs (that's the usual case), the antiresonance due to the mode coupling does not cause servo instability thank to the sufficient phase margin.

In conclusion, unfortunately we can't diagnalize the sensors and actuators using the natural modes because our assumption of the mode purity is not valid.
We can leave the pitch/yaw modes undiagnalized or just believe the oplevs as a relatively reliable reference of pitch and yaw and set the output matrix accordingly.

 

The figures will be rotated later.

Attachment 1: 161007_P.pdf
161007_P.pdf
Attachment 2: 161007_Y.pdf
161007_Y.pdf
  12539   Fri Oct 7 20:25:14 2016 KojiUpdateCDSPower-cycled c1psl and c1iool0

Found the MC autolocker kept failing, It turned out that c1iool0 and c1psl went bad and did not accept the epics commands.

Went to the rack and power cycled them. Burt resotred with the snapshot files at 5:07 today.

The PMC lock was restored, IMC was locked, WFS turned on, and WFS output offloaded to the bias sliders.

The PMC seemed highly misaligned, but I didn't bother myself to touch it this time.

  12538   Fri Oct 7 15:36:03 2016 SteveUpdatePEMthe lab temp is up

The IFO room temp is up

Quote:

The IFO room temp is up a bit and it is coming down. The out side temp is not really high.

 

Attachment 1: tempUp.png
tempUp.png
Attachment 2: PEM4d.png
PEM4d.png
  12537   Fri Oct 7 10:29:57 2016 SteveUpdateVACRGA removed

The RGA is removed for repaire. It's volume at atmophere and sealed.. P4 reading of 38 Torr is not correct.

 

Attachment 1: RGAremoved.png
RGAremoved.png
  12536   Thu Oct 6 15:42:51 2016 LydiaUpdateSUSOutput matrix diagonalization

Summary: At the 40m meeting yesterday, Eric Q. gave the suggestion that we accept the input matrix weirdness and adjust the output matrix by driving each coil individually so that it refers to the same degrees of freedom. After testing this strategy, I don't think it will work. 

Yesterday evening I tested this idea by driving one ITMY coil at a time, and measuring the response of each of the free swing modes at the drive frequency. I followed more or less the same procedure as the standard diagonalization: responses to each of the possible stimuli are compared to build a matrix, which is inverted to describe the responses given the stimuli. For the input matrix, the sensor readings are the responses and the free swing peaks are the stimuli. For the output matrix, the sensors transformed by the diagonalized input matrix as the responses of the dofs which are compared, and the drive frequency peak associated with a coil output is the stimulus. However, the normalization still happens to each dof independently, not to each coil independently. 

The output matrix I got had good agreement with the ITMY input matrix in the previous elog: for each dof/osem the elements had the same sign in both input and output matrices, so there are no positive feedback loops. The relative magnitude of the elements also corresponded well within rows of the input matrix. So the input and output matrices, while radically different from the ideal, were consistent with each other and referred to the same dof basis. So, I applied these new matrices (both input and output) to the damping loops to test whether this approach would work. 

drive-generated output matrix: 

      UL      UR      LR       LL      SD
pit    1.701  -0.188  -2.000  -0.111   0.452  
yaw    0.219  -1.424   0.356   2.000   0.370  
pos    1.260   1.097   0.740   0.903  -0.763  
sid    0.348   0.511   0.416   0.252   1.000  
but    0.988  -1.052   0.978  -0.981   0.060

However, when Gautam attempted to lock the Y arm, we noticed that this change significantly impacted alignment. The alignment biases were adjusted accordingly and the arm was locking. But when the dither was run, the lock was consistently destroyed. This indicates that the dither alignment signals pass through the SUS screen output matrix. If the output matrix pitch and yaw columns refer instead to the free swing eigenmodes, anything that uses the output matrix and attempts to align pitch and yaw will fail. So, the ITMY matrices were restored to their previous values: a close to ideal input matrix and naive output matrix. We could try to change everything that is affected by the output matrices to be independent of a transformation to the free swing dof basis, and then implement this strategy. But to me, that seems like an unneccessary amount of changes with unpredictable consequences in order to fix something that isn't really broken. The damping works fine, maybe even better, when the input matrix is set by the output matrix: we define pitch, for example, to be "The mode of motion produced by a signal to the coils proportional to the pitch row of the naieve output matrix," and the same for the other dofs. Then you can drive one of these "idealized" dofs at a time and measure the sensor responses to find the input matrix. (That is how the input matrix currently in use for ITMY was found, and it seems to work well.) 

 

  12535   Thu Oct 6 03:56:43 2016 ericqUpdateLSCRevival Attempt

[ericq, Gautam, Lydia]

We spent some time tonight trying to revive the PRFPMI. (Why PR instead of DR? Not having to deal with SRM alignment and potentially get a better idea of our best-case PRG). After the usual set up and warm up, we found ourselves unable to hold on to the PRMI while the arms flash. In the past, this was generally solved through clever trigger matrix manipulations, but this didn't really work tonight. We will meditate on the solution.

  12534   Wed Oct 5 19:43:13 2016 gautamSummaryGeneralVent review

This elog is meant to review some of the important changes made during the vent this summer - please add to this if I've forgotten something important. I will be adding this to the wiki page for a more permanent record shortly.


Vent objectives:

  1. Clean ITMX, ITMY, ETMX, ETMY
  2. Replace ETMX suspension cage, replace Al wire standoffs with Ruby (sapphire?) standoffs.
  3. Shorten Y arm length by 20mm
  4. Replace 40mm aperture baffles in ETM chambers with 50mm black glass baffles

Optics, OSEM and suspension status:

ITMX & ITMY

  • ITMX and ITMY did not have any magnets broken off during the vent - all five OSEM coils for both were removed and the optic EQ stopped for F.C. cleaning.
  • Both HR and AR faces were F.Ced, ~20mm dia area cleaned.
  • The coils were re-inserted in an orientation as close to the original (as judged from photos), and the shadow sensor outputs were made as close to half their open values as possible, although in the process of aligning the arms, this may have changed
  • OSEM filter existense was checked (to be updated)
  • Shadow sensor open values were recorded (to be updated)
  • Checked that tables were level before closing up
  • The UL OSEM on ITMY was swapped for a short OSEM while investigating glitchy shadow sensor outputs. This made no difference. However, the original OSEM wasn't replaced. Short OSEM was used as we only had spare short OSEMs. Serial number (S/N 228) and open voltage value have been recorded, wiki page will be updated. Does this have something to do with the input matrix diagonalization weirdness we have been seeing recently?
  • ITMX seems to be prone to getting stuck recently, reason unknown although I did notice the LL OSEM was kind of close to the magnet while inserting (but this magnet is not the one getting stuck, as we can see this clearly on the camera - the prime suspect is UL I believe)
  • OL beam centering on in vacuum steering optics checked before closing up

ETMY

  • UL, UR and LR magents broke off at various points, and so have been reglued
  • No standoff replacement was done
  • Re-suspension was done using newly arrived SOS wire
  • Original OSEMs were inserted, orientations have changed somewhat from their previous configuration as we did considerable experimentation with the B-R peak minimization for this optic
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • New wire suspension clamp made at machine shop is used, 5 in lb of torque used to tighten the clamp
  • HR face cleaned with F.C.
  • Optic + suspension towers air baked (separately) at 34C for curing of EP30
  • Checked that tables were level before closing up
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.
  • Suspension cage was moved towards ITMY by 19mm (measured using a metal spacer) by sliding along stop marking the position of the tower.

ETMX

  • Al wire standoffs <--> Ruby wire standoffs (this has changed the pitch frequency)
  • All magnets were knocked off at some point, but were successfully reglued
  • New SOS tower, new SOS wire, new wire clamp used
  • OSEM filter status, shadow sensor open voltage values to be updated.
  • OSEM orientation is close to horizontal for all 5 OSEMs
  • Table leveling was checked before closing up.
  • 40mm O.D. black glass baffle replaced with 50mm O.D. baffle.\

PRM

  • Some issues with the OSEMs were noticed, and were traced down to the Al foil caps covering the back of the (short) OSEMs, which are there to minimize the scattererd 1064nm light interfering with the shadow sensor, shorting one of the OSEMs
  • To mitigate this, all Al foil caps now have a thin piece of Kapton between foil and electrical contacts on rear of OSEM
  • No OSEMs were removed from the suspension cage during this process, we tried to be as gentle as possible and don't believe the shadow sensor values changed during this work, suggesting we didn't disturb the coils (PRM wasn't EQ stopped either)

SRM

  • The optic itself wasn't directly touched during the vent - but was EQ stopped as work was being done on ITMY
  • It initially was NOT EQ stopped, and the shift in table level caused by moving ITMY cage to the edge of the table for F.C. cleaning caused the optic to naturally drift onto the EQ stops, leading to some confusion as to what happened to the shadow sensor outputs
  • The problem was diagnosed and restoring ITMY to its original position made the OSEM signals come back to normal.

SR3

  • Was cleaned by drag wiping both front and back faces

SR2/PR2/PR3/BS/OMs

  • These optics were NOT intentionally touched during this vent
  • The alignment on the OMs was not checked before close-up
 

Other checks/changes

  • OL beams were checked on in-vacuum input and output steering mirrors to make sure none were close to clipping
  • Insides of viewport windows were checked for general cleanliness, given that we have found the outside of some of these to be rather dirty. Insides of viewports checked were deemed clean enough.
  • Steve has installed a new vacuum guage to provide a more realiable pressure readout. 
  • We forgot to investigate the weird behaviour of the AS beam that Yutaro and Koji identified in November. In any case, looks like the clipping of the AS beam is worse now. We will have to try and fix this using the PZT mounted OMs, and if not, we may have to consider venting again

Summary of characterization tasks to be done:

  1. Mode matching into the Y arm cavity given the arm length change
  2. HOM content in transmitted IR light from Y arm given the arm length change (Finesse models suggest that the 2f second order HOM resonance may have moved closer to the 00 resonance)
  3. Arm loss measurement
  4. Suspension diagonalization
  5. Check the Qs of the optics eigenmodes - should indicate if any of our magnets, reglued or otherwise, are a little loose
  12533   Wed Oct 5 19:10:04 2016 gautamUpdateGeneralArm loss measurement review

[ericq,gautam]

There are multiple methods by which the arm loss can be measured, including, but not limited to:

  1. Cavity ringdown measurement
  2. Monitoring IR arm transmission using ALS to scan the arm through multiple FSRs
  3. Monitoring the reflected light from the ITM with and without a cavity (Johannes has posted the algebra here)

We found that the second method is extremely sensitive to errors in the ITM transmissivity. The first method was not an option for a while because the AOM (which serves as a fast shutter to cut the light to the cavity and thereby allow measurement of the cavity ringdown) was not installed. Johannes and Shubham have re-installed this so we may want to consider this method.

Most of the recent efforts have relied on the 3rd method, which itself is susceptible to many problems. As Yutaro found, there is something weird going on with ASDC which makes it perhaps not so reliable a sensor for this measurement (unfortunately, no one remembered to follow up on this during the vent, something we may come to regret...sad). He performed some checks and found that for the Y arm, POY is a suitable alternative sensor. However, the whitening gain was at 0dB for the measurements that Johannes recently performed (Yutaro does not mention what whitening gain he used, but presumably it was not 0). As a result, the standard deviation during the 10s averaging was such that the locked and misaligned readings had their 'fuzz' overlapping significantly. The situation is worse for POX DC - today, Eric checked that the POX DC and POY DC channels are indeed reporting what they claim, but we found little to no change in the POX DC level while misaligning the ITM - even after cranking the whitening gain up to 40!

Eric then suggested deriving ASDC from the AS110 photodiode, where there is more light. This increased the SNR significantly - in a 10s averaging window, the fuzz is now about 10 ADC counts out of ~1500 (~<1%) as opposed to ~2counts out of 30 previously. We also set the gains of POX DC, POY DC and ASDC to 1 (they were 0.001,0.001 and 0.5 respectively, for reasons unknown).

I ran a quick measurement of the X arm loss with the new ASDC configuration, and got a number of 80 +/- 10 ppm (7 datapoints), which is wildly different from the ~250ppm number I got from last night's measurement with 70 datapoints. I was simultaneously recording the POX DC value, which yielded 40 +/- 10 ppm.

We also discovered another possible problem today - the spot on the AS camera has been looking rather square (clearly not round) since, I presume, closing up and realigning everything. By looking at the beam near the viewport on the AS table for various configurations of the ITM, we were able to confirm that whatever is causing this distortion is in the vacuum. By misaligning the ITM, we are able to recover a nice round spot on the AS camera. But after running the dither align script, we revert to this weirdly distorted state. While closing up, no checks were done to see how well centered we are on the OMs, and moreover, the DRMI has been locked since the vent I believe. It is not clear how much of an impact this will have on locking the IFO (we will know more after tonight). There is also the possibility of using the PZT mounted OMs to mitigate this problem, which would be ideal.


Long story short -

  1. Some more thought needs to be put into the arm loss measurement. If we are successful in locking the IFO, the PRG would be a good indicator of the average arm loss.
  2. There is some clipping, in vacuum, of the AS beam. It may be that we can fix this without venting, to be investigated.
 

GV Edit 8 Oct 2016: Going through some old elogs, I came across this useful reference for loss measurement. It doesn't talk about the reflection method (Method 3 in the list at the top of this elog), but suggests that cavity ringdown with the Trans PD yields the most precise numbers, and also allows for measuring TITM

ELOG V3.1.3-